OpenStudy (jason613):

Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2

Mehek (mehek14):

Point slope form: \(\bf \large{y-y_1 = m(x-x_1)}\) m = slope x1 and y1 are from the point you have, so all you need to do is plug in the values.

Mehek (mehek14):

@jason613 Do you understand?

OpenStudy (jason613):

no

Mehek (mehek14):

Alright, the point-slope form is \(\bf \large y-y_1=m(x-x_1)\) Do you understand that?

OpenStudy (joannablackwelder):

Do you have options given for the answer? Multiple choice?